A car travels 40 miles an hour and a plane travels 10 miles in a minute. How far will the car travel while the plane travels 450 miles?
Answer by mananth(16946)

Plane 10 miles a minute=10*60=600 mph
t(plane)=d/r
450/600 = 3/4 hours
D(car)= 40*3/4= 30 miles